OpenStudy (dinamix):

hmm i want proof for i^2 = -1 why , how did get that , just for enjoyment so anyone help me ?

OpenStudy (anonymous):

its an imaginary number

OpenStudy (anonymous):

that's whats it called in math terms

OpenStudy (anonymous):

that's what i stand for too.

OpenStudy (dinamix):

yup i know but why and how we get this number

OpenStudy (dinamix):

@ganeshie8

OpenStudy (anonymous):

you cant get -1 from a exponent so what they do is that they just put an "i"

OpenStudy (amistre64):

does: x^2 + 1 = 0 have a solution?

OpenStudy (dinamix):

in R no but in C have solution ;p

OpenStudy (ikram002p):

lets talk about complex invention :3

OpenStudy (amistre64):

x^2 = -1 is the solution; now can we come up with a consistent mathical philosophy with this concept? if so then it is valid, otherwise it is just a whimsical consequence.

OpenStudy (dinamix):

For this reason I opened subject Mr.amistre 64

OpenStudy (amistre64):

you cannot 'prove' that i^2 = -1, it is an axiom ; or definition that creates a consistent mathical philosophy.

OpenStudy (dinamix):

when will be mathical philosophy and how @amistre64 ;p ?

OpenStudy (ikram002p):

:D its like saying prove that e=something